OpenStudy (anonymous):

A circle has its center at (1, 4) and a radius of 2 units. What is the equation of the circle? Answer (x + 1)2 + (y - 4)2 = 4 (x - 1)2 + (y - 4)2 = 4 (x - 1)2 + (y - 4)2 = 2 (x + 2)2 + (y + 4)2 = 2

OpenStudy (amistre64):

the copy paste makes for a rather large question ....

OpenStudy (amistre64):

the equation of a circle; with a given center (cx,cy) and a given radius "r" is of the form: \[(x-c_x)^2+(y-c_y)^2=r^2\]

OpenStudy (anonymous):

(x - 1)2 + (y - 4)2 = 4